OpenStudy (anonymous):

Can someone please show how to solve this inequality? -3x + 4

OpenStudy (amistre64):

already showed you; only thing left to do is to DO it for you....

OpenStudy (amistre64):

what do you get when you add 3x to both sides?

OpenStudy (anonymous):

When I mean SHOW I'm referring to a step by step interpretation of how to solve the problem. not directions.

OpenStudy (anonymous):

Im confused in which sides

OpenStudy (amistre64):

both sides: -3x+(3x) +4 </ x +(3x) +8

OpenStudy (amistre64):

there is a left side and a right side to the </ sign right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

intsead of an (=) betweent hem; you have the </ sign.... thats the only diff

OpenStudy (amistre64):

how would you solve: -3x +4 = x +8 ??

OpenStudy (anonymous):

add 3 to both sides?

OpenStudy (amistre64):

add (3x) to both sides would be more helpful :) but ill assume thats a typo..

OpenStudy (amistre64):

-3x+3x +4 = x+3x +8 0+4 = 4x+8 4 = 4x+8 now what?

OpenStudy (anonymous):

divide by 4

OpenStudy (amistre64):

divide by for is possible; but there is an easier step first... we can -8 from each side to get that 4x by itself

OpenStudy (amistre64):

4-8 = 4x +8-8 -4 = 4x + 0 -4 = 4x ; now we can easily /4 -4/4 = 4x/4 -1 = 1x -1=x

OpenStudy (amistre64):

since we dint mess with multiplying or dividing the 'x' by a negative number; we can use the same sign -1 </ x

OpenStudy (amistre64):

draw a vertical line thru x =-1 and since -1 </ 0 is true; the origin is included in the area to be shaded

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

youre welcome :)